# End-to-End Large Portfolio Optimization for Variance Minimization with Neural Networks through Covariance Cleaning

### v2

@@ 4.4.2 Frictionless Interpolation Backtesting
To assess the performance in a statistically robust way, we performed a battery of 1,0001,0001 , 000 independent backtests. Specifically, every simulation consists of approximately 555 years of data, starting from a random day $t$ within [2000-01-01, 2019-12-31], and performing 250250250 portfolio rebalancing every 555 trading days. Our model will use the last available calibration obtained the year before the trading day so that no data leakage occurs. The list of assets is randomly sampled from the investable universe (see Sec. 4.1) on the first day of the simulation, and it is kept fixed if possible for the entire backtest. If some of the assets are excluded from the investable universe during the backtest, they are removed from the portfolio, and another random asset is added to the portfolio so that the number of assets remains constant.
Method $\mathcal{L}(\bm{w},\bm{\Sigma}_{\mathrm{out}})$ Volatility Mean Return Sharpe Ratio $\mathbf{n_{\mathrm{eff}}}$ Turnover Leverage
@@
Table 1: Performance comparison of frictionless portfolio backtests for both unconstrained (long-short) and constrained (long-only) settings. The table reports averages over 1,0001{,}0001 , 000 independent simulations of 250250250 rebalancings (every 5 trading days) on portfolios of $n=300$ stocks randomly sampled during [2000-01-01,2019-12-31], without transaction costs. Values are sorted by Sharpe ratio, and bold entries indicate the statistically significant top rank (percentile bootstrap, $p<10^{-6}$). Portfolio volatility, mean return, and Sharpe ratio are annualized; $n_{\mathrm{eff}}$ is computed via the inverse Herfindahl index [78]. The loss function $\mathcal{L}(\bm{w},\bm{\Sigma}_{\mathrm{out}})$ is defined in Eq.(13) and refers to the short-term variance ($\Delta t_{\textrm{out}}=5$ days).
[…]
@@
The other two columns in Tab. 1 report the average number of effective assets in the portfolio. This metric is dominated by the AO estimator, which has an average of 33.233.233.2 effective assets, while the others range between 13.313.313.3 and 17.617.617.6, which are less than half of the number of assets in the portfolio.
[…]
@@ 4.4.3 High Realistic Extrapolation Backtesting
The simulator returns, for every calendar date, the full path of $\text{NLV}_{t}$, cash, market values, and margin statistics, thus reflecting the realistic interaction between turnover, leverage due to slippage, and transaction costs. The simulation starts with a cash balance of 1,$0$, which is a reasonable amount for a small hedge fund portfolio, and it is sufficiently low to avoid the need to account for market impact on the most liquid US stocks [82]. Our portfolio is rebalanced every 555 trading days, and the investable asset universe is taken from the procedure described in Sec. 4.1, with the only difference that we select the top $n=1,000$ for marketcap, defined from the previous day’s close price. In this case, we are attempting to extrapolate the NN to a larger number of assets, given that the calibration was done on a maximum of 350350350 assets. Furthermore, $q=n/\Delta t_{\text{in}}\approx 0.83$, which is substantially larger than the training range of $q\in[0.04,0.29]$.
@@
Method Short Var. Sharpe Ratio Sortino Ratio Mean Return Volatility Turnover $\mathbf{n_{\mathrm{eff}}}$
NN 9.0% 1.058 1.268 12.6% 11.9% 57.0% 46.0
AO 9.7% 0.942 1.155 11.7% 12.5% 18.0% 47.8
PM 9.8% 0.870 1.077 10.8% 12.5% 23.0% 25.7
QIS 9.8% 0.848 1.053 10.6% 12.6% 23.0% 20.1
MLE 9.8% 0.847 1.052 10.6% 12.5% 25.0% 18.8
ERB 20.2% 0.533 0.710 13.0% 24.5% 6.0% 580.2
MCW 15.7% 0.484 0.615 9.4% 19.5% 7.0% 146.6
@@
Table 2: Performance of the long-only portfolio with $n=1000$ top capitalization stocks, rebalanced every 5 trading days, from January 2000 to December 2024. The simulations use a realistic trading simulator that accounts for transaction costs and slippage. Ratios, returns, and volatility are annualized. The turnover metric is computed without accounting for price drift and serves only as a reference to the model’s stability, $n_{\text{eff}}$ is computed by using the inverse Herfindahl index [78]. The first column indicates the average short-term portfolio variance ($\Delta t_{\textrm{out}}=5$) consistently with the NN loss function (Eq. (13)). Bold values indicate the best-ranked model.
@@
In Tab. 2, we can observe the performance metrics of the portfolio across different strategies. Notably, when considering the realized variance over the five-day OOS window (i.e., the first column of the table), our NN model again achieves the lowest variance, a result that is fully consistent with its calibration objective. The NN strategy outperforms the others in terms of Sharpe ratio, Sortino ratio, volatility, and reaches the second-best mean return. The second-best Sharpe ratio is achieved by the AO estimator. All the other methods, including the MLE covariance estimator, have similar performances in terms of Sharpe ratio, Sortino ratio, and volatility. The univariate methods, such as ERB and MCW, perform significantly worse than the multivariate methods, with a Sharpe ratio of 0.6780.6780.678 and 0.4830.4830.483, respectively. In fact, for the ERB strategy, which allocates weights by inverting individual variances in a univariate fashion, the diversification indicator $n_{\text{eff}}$ reaches 580 out of 1,000, reflecting a relatively uniform allocation. As detailed in Sec. 4.1, stocks with exceptionally low past volatility were excluded, thereby preserving the multivariate nature of the optimization problem.
@@
Figure 6: Long-only portfolio performances of the top 1,000 most capitalized stocks in the universe backtested with the high realistic simulator. Top left: NLV of the portfolio as a function of time. Top right: Annualized volatility of the portfolio as a function of time over a yearly rolling window. Bottom: MDD of the portfolio by year.
@@ 5 Discussions
For future work, we see four areas for further improvement. First, one can replace the rotation-invariant eigenvalue filter with a context-dependent cleaning that adapts to evolving market regimes. Second, denoising can be extended to eigenvectors, a task that hierarchical clustering methods already accomplish successfully [23, 20, 21]. Third, the modules could be coupled so that lag-transformation, eigenvalue filtering, and volatility scaling interact dynamically rather than operating independently. Finally, we plan to generalize our framework to the full mean-variance problem by integrating return estimation with covariance learning in a single end-to-end network, possibly including differentiable optimization layers to handle non-closed-form objectives. This extension would offer a more exhaustive solution to portfolio selection.
Author Contributions
C.B. conceived the study, designed the NN architecture, developed the code implementation, conducted the numerical simulations, and drafted the manuscript. E.M. contributed to the code development, carried out the exploratory simulations, data analysis, prepared the initial draft of the manuscript, and participated in the revision. R.N.M. reviewed the methodology and results, participated in discussions to refine the analysis, and contributed to the manuscript revisions.
Acknowledgments
This work was performed using HPC resources from the “Mésocentre” computing center of CentraleSupélec and École Normale Supérieure Paris-Saclay supported by CNRS and Région Île-de-France (http://mesocentre.centralesupelec.fr/). E.M. acknowledges a fellowship funded by PNRR for the PhD DOT1608375 in Sistemi complessi per le scienze fisiche, socio-economiche e della vita of Catania University.

### v3

@@ 4.4.2 Frictionless Interpolation Backtesting
To assess the performance in a statistically robust way, we performed a battery of $1,000$ independent backtests. Specifically, every simulation consists of approximately $5$ years of data, starting from a random day $t$ within [2000-01-01, 2019-12-31], and performing $250$ portfolio rebalancing every $5$ trading days. Our model will use the last available calibration obtained the year before the trading day so that no data leakage occurs. The list of assets is randomly sampled from the investable universe (see Sec. 4.1) on the first day of the simulation, and it is kept fixed if possible for the entire backtest. If some of the assets are excluded from the investable universe during the backtest, they are removed from the portfolio, and another random asset is added to the portfolio so that the number of assets remains constant.
Method $\mathcal{L}(\bm{w},\bm{\Sigma}_{\mathrm{out}})$ Vol. Return Sharpe $\mathbf{n_{\mathrm{eff}}}$ Turn. GL
@@
Table 1: Performance comparison of frictionless portfolio backtests for both unconstrained (long-short) and constrained (long-only) settings. The table reports averages over 1,0001{,}000 independent simulations of $250$ rebalancings (every 5 trading days) on portfolios of $n=300$ stocks randomly sampled during [2000-01-01,2019-12-31], without transaction costs. Values are sorted by Sharpe ratio, and bold entries indicate the statistically significant top rank (percentile bootstrap, $p<10^{-6}$). Portfolio volatility, mean return, and Sharpe ratio are annualized; $n_{\mathrm{eff}}$ is computed via the inverse Herfindahl index [78]. The loss function $\mathcal{L}(\bm{w},\bm{\Sigma}_{\mathrm{out}})$ is defined in Eq.(13) and refers to the short-term variance ($\Delta t_{\textrm{out}}=5$ days). Turnover is measured excluding price drift, and gross leverage is one by definition under long-only constraints. Abbreviations: NN = Neural Network estimator; MLE = Maximum-Likelihood (sample) Estimator; QIS = Quadratic-Inverse Shrinkage; LS = Linear Shrinkage; PM = Power Mapping; AO = Average Oracle; Vol. = annualized Volatility; Turn. = Turnover; GL = Gross Leverage; $n_{\mathrm{eff}}$ = Effective Number of assets (inverse Herfindahl index).
In Tab. 1, we report the average performances of the models tested on the same set of assets. The first column displays the short-term portfolio variance, which coincides with the neural network loss function defined in Eq. (13). As expected, our trained model outperforms all other estimators in the metric that it optimizes. The unconstrained (long–short) results show that the NN achieves the highest Sharpe ratio (1.01) with an annualized volatility of $10.9\%$ and a mean return of $11.0\%$. The QIS estimator attains the second‐best Sharpe ratio ($0.94$) but does so with a notably high gross leverage of $3.81$, compared with the NN leverage of 3.07. The AO method ranks third in Sharpe ratio ($0.91$) yet operates with the lowest leverage ($2.13$) and minimal turnover ($16.3\%$). These performance gains for the NN and QIS come at the cost of trading intensity. The network rebalances more than its entire notional every five trading days (turnover $112.0\%$), and QIS rebalances $44.0\%$ of its portfolio at each rebalance. Such high turnover and leverage would incur significant financing and borrowing costs not captured in this frictionless experiment.
Once short-selling is forbidden (bottom of Tab. 1), most of the models perform similarly. This is a known behavior of covariance cleaning for portfolio optimization [63, 39]. The annualized volatility is around $14.2\%$ for all the models, including the MLE covariance estimator, which is the non-cleaned covariance matrix. The AO estimator performs slightly worse, reaching $14.6\%$ annualized volatility. The NN instead reaches $13.5\%$ annualized volatility, which is the lowest among all the models by a substantial margin. In terms of the Sharpe ratio, the NN reaches a $0.79$, which is the highest among all the models, while the AO estimator reaches the second-best Sharpe ratio of $0.74$, while all the other methods range around $0.72$. We want to stress that checking the Sharpe ratio is very important when measuring the performance of a portfolio in a large, uncontrolled setting. In fact, it might be possible that the data cleaning included stocks that must be excluded, for example, stocks that are about to be delisted or merged or stocks that are not liquid enough. In those cases, the asset might have a very low volatility and a very low return. A portfolio that includes those assets will have a very low volatility, but it will also have a very low return, leading to a very low Sharpe ratio. The objective of a GMV portfolio is, in fact, to minimize the volatility of investing in risky assets.
@@
The other two columns in Tab. 1 report the average number of effective assets in the portfolio. This metric is dominated by the AO estimator, which has an average of $33.2$ effective assets, while the others range between $13.3$ and $17.6$, which are less than half of the number of assets in the portfolio.
Finally, the average turnover reported in the last column measures the fraction of assets that are changed in the portfolio at every rebalancing. The AO estimator has the lowest turnover ($10\%$) while the NN has a turnover of $47\%$, which is the highest among all the models. This is due to the fact that the NN is trained to be reactive and to focus on the most recent past, while the other models are more stable and less sensitive to recent changes. This behavior is a double-edged sword; on one side, having a more reactive portfolio allows one to adapt to the market conditions, but on the other side, it leads to a higher turnover and, therefore, higher transaction costs. For this reason, in the next section, we will evaluate the performance of the NN in a more realistic setting.
@@ 4.4.3 High Realistic Extrapolation Backtesting
The simulator returns, for every calendar date, the full path of $\text{NLV}_{t}$, cash, market values, and margin statistics, thus reflecting the realistic interaction between turnover, leverage due to slippage, and transaction costs. The simulation starts with a cash balance of 1,000,000 \$, which is a reasonable amount for a small hedge fund portfolio, and it is sufficiently low to avoid the need to account for market impact on the most liquid US stocks [67]. Our portfolio is rebalanced every $5$ trading days, and the investable asset universe is taken from the procedure described in Sec. 4.1, with the only difference that we select the top $n=1,000$ for marketcap, defined from the previous day’s close price. In this case, we are attempting to extrapolate the NN to a larger number of assets, given that the calibration was done on a maximum of $350$ assets. Furthermore, $q:=n/\Delta t_{\text{in}}\approx 0.83$, which is substantially larger than the training range of $q\in[0.04,0.29]$.
Evaluating the trained model on a larger universe highlights the practical value of the dimension agnostic design. Because the parameters are shared across assets and eigenvalue ranks, a network calibrated on moderate cross sections can be deployed at inference on substantially larger panels without retraining, which reduces the training cost and simplifies model maintenance when the investable universe changes. This extrapolation test also provides evidence that the learned spectral transformation captures a stable, transferable regularization mechanism rather than idiosyncrasies of a specific panel.
@@
Method $\bm{\mathcal{V}_{5}}$ Sharpe Sortino Mean Vol. VaR5% CVaR5% Turn. $\mathbf{n_{\mathrm{eff}}}$
NN 9.0% 1.058 1.268 12.6% 11.9% -1.05% -1.76% 57.0% 46.0
AO 9.7% 0.942 1.155 11.7% 12.5% -1.12% -1.85% 18.0% 47.8
PM 9.8% 0.870 1.077 10.8% 12.5% -1.13% -1.84% 23.0% 25.7
QIS 9.8% 0.848 1.053 10.6% 12.6% -1.10% -1.84% 23.0% 20.1
DCC 10% 0.930 1.177 11.5% 12.4% -1.1% -1.80% 54% 17.0
MLE 9.8% 0.847 1.052 10.6% 12.5% -1.10% -1.84% 25.0% 18.8
ERB 20.2% 0.533 0.710 13.0% 24.5% -2.41% -3.59% 6.0% 580.2
MCW 15.7% 0.484 0.615 9.4% 19.5% -1.88% -2.95% 7.0% 146.6
@@
Table 2: Performance of the long-only portfolio with $n=1000$ top capitalization stocks, rebalanced every 5 trading days, from January 2000 to December 2024. The simulations use a realistic trading simulator that accounts for transaction costs and slippage. Ratios, returns, and volatility are annualized. The turnover metric is computed without accounting for price drift and serves only as a reference to the model’s stability, $n_{\text{eff}}$ is computed by using the inverse Herfindahl index [78]. The first column indicates the average short-term portfolio variance ($\Delta t_{\textrm{out}}=5$) consistently with the NN loss function (Eq. (13)). In addition to variance-based metrics, we report tail-risk measures (VaR and CVaR) computed from the realised portfolio return series at the 5% left tail. Values are sorted by Sharpe ratio, and bold values indicate the best-ranked model. Abbreviations: NN = Neural Network estimator; MLE = Maximum-Likelihood (sample) Estimator; QIS = Quadratic-Inverse Shrinkage; PM = Power Mapping; AO = Average Oracle; DCC = Dynamic Conditional Correlation; ERB = Equal-Risk Budget; MCW = Market-Capitalization Weighted; VaR5% = 5% Value-at-Risk; CVaR5% = 5% Conditional Value-at-Risk; Vol. = annualized Volatility; Turn. = Turnover; $n_{\mathrm{eff}}$ = Effective Number of assets (inverse Herfindahl index).
@@
In Tab. 2, we can observe the performance metrics of the portfolio across different strategies. Notably, when considering the realized variance over the five-day OOS window (i.e., the first column of the table), our NN model again achieves the lowest variance, a result that is fully consistent with its calibration objective. This risk reduction is corroborated by the quantile-based tail measures reported in the table, as both Value-at-Risk (VaR) and Conditional Value-at-Risk (CVaR) are also improved by the NN. Furtheremore, the NN strategy outperforms the others also in terms of Sharpe ratio, Sortino ratio, and reaches the second-best mean return. The second-best Sharpe ratio is achieved by the AO estimator. The DCC benchmark ranks close behind (Sharpe $0.93$) but remains below both NN and AO. All the other methods, including the MLE covariance estimator, have similar performances in terms of Sharpe ratio, Sortino ratio, and volatility. The univariate methods, such as ERB and MCW, perform significantly worse than the multivariate methods, with a Sharpe ratio of $0.678$ and $0.483$, respectively. In fact, for the ERB strategy, which allocates weights by inverting individual variances in a univariate fashion, the diversification indicator $n_{\text{eff}}$ reaches 580 out of 1,000, reflecting a relatively uniform allocation. As detailed in Sec. 4.1, stocks with exceptionally low past volatility were excluded, thereby preserving the multivariate nature of the optimization problem.
[…]
@@
Figure 6: Long-only portfolio performances of the top 1,000 most capitalized stocks in the universe backtested with the high realistic simulator. Top left: NLV of the portfolio as a function of time. Top right: Annualized volatility of the portfolio as a function of time over a yearly rolling window. Bottom: MDD of the portfolio by year. Abbreviations: NN = Neural Network estimator; MLE = Maximum-Likelihood (sample) Estimator; AO = Average Oracle; MCW = Market-Capitalization Weighted; NLV = Net Liquidation Value; MDD = Maximum Drawdown.
Finally, we address whether there exists an optimal aspect ratio $q_{\textrm{in}}:=n/\Delta t_{\mathrm{in}}$ for the scalable model. In our setup, $\Delta t_{\mathrm{in}}$ is fixed, hence varying the universe size $n$ directly sweeps $q_{\textrm{in}}$. Tab. 3 shows that, over the tested range, performance improves monotonically as $n$ increases (equivalently, as $q_{\textrm{in}}$ increases), with the best results attained at $n=1000$ (corresponding to $q_{\textrm{in}}\simeq 0.83$). We therefore do not observe an optimal-$q_{\textrm{in}}$ window in the interior in these experiments. This trend is consistent with the fact that larger universes offer more diversification opportunities, while the main mechanism that would otherwise degrade GMV performance at large $q_{\textrm{in}}$ is the growth of sampling noise, which our denoising procedure precisely mitigates.
$\mathbf{n}$ $\bm{\mathcal{V}_{5}}$ Sharpe Sortino Mean Vol. VaR5% CVaR5% Turn. $\mathbf{n_{\mathrm{eff}}/n}$
21 12.6% 0.513 0.669 8.0% 15.5% -1.44% -2.31% 18.3% 56.6%
55 11.4% 0.575 0.752 8.1% 14.2% -1.29% -2.07% 26.2% 19.4%
144 10.6% 0.680 0.847 9.1% 13.4% -1.20% -1.99% 29.2% 12.1%
379 9.9% 0.921 1.160 11.6% 12.6% -1.15% -1.85% 32.4% 5.8%
1000 9.0% 1.058 1.268 12.6% 11.9% -1.05% -1.76% 57.1% 4.6%
Table 3: Performance of the NN long-only portfolio as a function of the investment universe size $n$ (top-capitalization stocks), rebalanced every 5 trading days, from January 2000 to December 2024, using a realistic trading simulator that accounts for transaction costs and slippage. Ratios, returns, and volatility are annualized. The turnover metric is computed without accounting for price drift and serves only as a reference to the model’s stability. The last column reports the diversification ratio $n_{\mathrm{eff}}/n$ (in percent), with $n_{\mathrm{eff}}$ computed via the inverse Herfindahl index [78]. The first column reports the average short-term portfolio variance at $\Delta t_{\mathrm{out}}=5$, consistently with the NN loss function (Eq. (13)). Abbreviations: VaR5% = 5% Value-at-Risk; CVaR5% = 5% Conditional Value-at-Risk; Vol. = annualized Volatility; Turn. = Turnover; $n_{\mathrm{eff}}$ = Effective Number of assets (inverse Herfindahl index).
@@ 5 Discussions
Despite the empirical improvements documented above, our current architecture remains constrained by the same structural assumptions that underlie RIEs. In particular, the eigenvalue-cleaning module acts only on the empirical spectrum and preserves the empirical eigenvectors, so the method cannot explicitly exploit, nor denoise, the time-varying information carried by eigenvector dynamics (e.g., evolving sector structure or changing factor exposures). As a consequence, the model should not be interpreted as a fully dynamic covariance forecaster or as a mechanism that detects regime shifts in a principled sense. Rather, it provides a data-driven regularisation of finite-sample noise within a static spectral framework, and it can only react to non-stationarity indirectly through the learned lag reweighting and clipping applied to the fixed-length input window. A second practical limitation is that the lag-transformation block allocates separate parameters to each look-back lag, which fixes calibration window length at training time and prevents a genuinely adaptive horizon without architectural changes.
These limitations motivate several concrete extensions. First, one can move beyond rotational invariance by replacing the purely spectrum-based filter with a context-dependent cleaning that is explicitly conditioned on market state variables or on temporal features, thereby enabling genuinely time-varying covariance representations. Second, denoising can be extended to eigenvectors, for instance by incorporating clustering-based structure that modifies both eigenvalues and eigenvectors, as suggested by the empirical success of hierarchical filters. Third, the modules could be coupled so that lag-transformation, eigenvalue filtering, and volatility scaling interact dynamically rather than operating as largely separate components. Finally, we plan to generalize the framework beyond GMV by integrating return estimation and/or differentiable constrained optimization layers, so that the training objective can directly reflect practical constraints (e.g., long-only mandates or turnover control) instead of relying on external post-processing.